OpenStudy (anonymous):

One day, a person went to horse racing area, Instead of counting the number of human and horses, he instead counted 74 heads and 196 legs. Yet he knew the number of humans and horses there. How did he do it, and how many humans and horses are there?

OpenStudy (anonymous):

humans have two legs horses have four legs but humans and horses only have one head

OpenStudy (anonymous):

total of legs 196 let n=#humans let m=#horses

OpenStudy (anonymous):

n(2)+m(4)=196

OpenStudy (anonymous):

but n+m=74

OpenStudy (anonymous):

rearrange for n n=74-m

OpenStudy (anonymous):

ok thanks man :D

OpenStudy (anonymous):

o u know what to do from there?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

the first few steps are the hardest